OpenStudy (anonymous):

[2/3 y -10] [2/3 y-10] i got 2/3 ysquared-100 is this correct

OpenStudy (anonymous):

[2/3 y -10][2/3 y +10]

OpenStudy (tkhunny):

Not very close. (a-b)*(a-b) = a^2 - 2ab + b^2 You have (2/3 y)^2 - 2(2/3 y)(10) + 10^2 = \(\dfrac{4}{9}y^{2} - \dfrac{40}{3}y + 100\)

OpenStudy (tkhunny):

(a+b)(a-b) = a^2 - b^2 You have (2/3 y)^2 - 10^2 = \(\dfrac{4}{9}y^{2} - 100\) You must get the coefficient, too, not just the variable.

OpenStudy (anonymous):

i was on the right track just must of some how added instead of multpily